OpenStudy (ltrout):

What is the equation of y = x^3 with the given transformations? - Vertical stretch by a factor of 3 - Horizontal shift 4 units to the right - Vertical shift 3 units down

OpenStudy (ltrout):

So far this is what I have: y = 3(x-4)^3. Stretched vertically by 3 and 4 units to the right

OpenStudy (anonymous):

yeah thats right now just subtract by 3 for a vertical shift down

OpenStudy (ltrout):

So it would be y = 3(x-4)^3 - 3?

OpenStudy (anonymous):

\[y = 3(x-4)^3 -3\] yep lol

OpenStudy (ltrout):

Okay lol figured. Was overthinking it :P. Thanks :)

OpenStudy (anonymous):

good looks for the medal

OpenStudy (anonymous):

a vertical stretch (multiply the function by 3): y = 3x^3 horizontal shift of 4 units to the right (subtract 4 units from the x): y = 3(x - 4)^3 Vertical shift 3 units down (subtract 3 units from the function): y = 3(x - 4)^3 - 3

OpenStudy (ltrout):

@tburn25 or @Marygirl21 can you guys help on one more?

OpenStudy (anonymous):

sure